You are given the following: Bond 1 Bond 2 Coupon rate of 4%, payable semi-annually Coupon rate of 8%, payable quarterly Effe0.0500 0.075 0.0975 0.1200 0.1025
Answer #1

OPTION E

From Bond 1:
Semiannual compounded yield=(1.05^(1/2)-1)*2=4.9390153%

Y=4%*100/4.9390153%*(1-1/(1+4.9390153%/2)^40)+100/(1+4.9390153%/2)^40=88.153300

From Bond 2:
Y=8%*100/r*(1-1/(1+r/4)^40)+100/(1+r/4)^40
=>8%*100/r*(1-1/(1+r/4)^40)+100/(1+r/4)^40=88.153300
=>r=9.87803066%

=>X=(1+9.87803066%/4)^4-1=10.25000002%